OpenStudy (ashley1nonly):

When I solve a problem and both have the same line. Does that mean that all points on the line is a solution?

OpenStudy (ashley1nonly):

\[\frac{ 1 }{ 2 }x-\frac{ 1 }{ 3 }y=1\] \[y=\frac{ 3 }{ 2 }x-3\]

OpenStudy (ashley1nonly):

\[-2x+\frac{ 4 }{ 3 }y=4\]

OpenStudy (ashley1nonly):

\[y=\frac{ 3 }{ 2 }x-3\]

OpenStudy (ashley1nonly):

The both have the same line so all points on the line are the solution.

OpenStudy (triciaal):

yes many solutions

OpenStudy (ashley1nonly):

Thank you

OpenStudy (triciaal):

you are welcome